OpenStudy (anonymous):

Solve –2(x – 1) = 10 Answers: x = –4 x = 4 x = 6 x = –6

OpenStudy (anonymous):

x = -4

OpenStudy (anonymous):

Thank you! how did you get this?

OpenStudy (anonymous):

i'll show you -2(x-1) = 10 -2x +2 = 10 (distribution) -2x = 8 (plus 2 or "move" 2 to the right of "=") x = -4 (divided by -2)